OpenStudy (anonymous):

solve the absolute value inequality and graph the solution on the real number line. |2x-3|greater than or equal to 5

OpenStudy (anonymous):

for absolute value , you need to squared the inequation |2x-3|≥5 4x^2 -12x+9≥25 4x^2 -12x-16≥0 divide all by 4 -> x^2 -3x-4≥0 (x-4)(x+1) ≥ 0 x = 4 x=-1 x ≤ -1 or x ≥ 4

OpenStudy (anonymous):

oh no do not square anything. \[|x|\geq p\] means \[x\geq p\] or \[x\leq -p\] so solve \[2x-3\geq5\] \[2x\geq 8\] \[x\geq4\] or \[2x-3 \leq-5\] \[2x \leq-2\] \[x\leq-1\]